OpenStudy (anonymous):

Use graphs and tables to find the limit and identify any vertical asymptotes of limit of 1 divided by the quantity x minus 9 as x approaches 9 from the left. The answers are : negative infinity; x = -9 1; no vertical asymptotes infinity; x = -9 negative infinity; x = 9

OpenStudy (anonymous):

I have no clue where to start on this. Help @freckles @xapproachesinfinity

OpenStudy (freckles):

\[\lim_{x \rightarrow 9^-} \frac{1}{x-9}\]

OpenStudy (freckles):

do you know what the graph of 1/(x-9) looks like?

OpenStudy (anonymous):

Yes that's the problem. No I don't.

OpenStudy (freckles):

and if you don't ... the graph of y=1/x is a pretty basic graph you will just move y=1/x to the right 9 units

OpenStudy (freckles):

 this graph is 1/x